# Multi-hop Reasoning via Early Knowledge Alignment

### v1

@@ 1 Introduction
Iterative RAG systems(Search-R1; R1-Searcher) are often optimized by Reinforcement Learning (RL)(PPO; shao2024deepseekmath), offering a principled approach to learn effective retrieval and reasoning strategies. RL-based RAG frameworks treat the retrieval and generation process as a sequential decision-making problem, where agents learn to search for information and generate responses to maximize cumulative rewards based on answer accuracy and efficiency metrics. The success of RL training heavily depends on the quality of the exploitation and the exploration efficiency during the learning process. Recent studies on entropy(wang2025beyond; cui2025entropy) show that entropy measurement is a good signal for this exploitation and exploration balance, which is important because the exploitation of retrieved information and exploration in the retrieval set control the whole reasoning process. Poor initial reasoning steps in exploration can lead to compounding errors throughout the iterative process.
@@
  • Analysis from an Entropy Perspective. We analyze the training dynamics of Group Relative Policy Optimization (GRPO)(shao2024deepseekmath) in iterative RAG from an entropy perspective and show that with lower entropy in the training phase, instead of insufficient exploration, our approach leads to more efficient exploration strategies focusing on the retrieval set, faster roads to the answer during RL training compared to traditional approaches that start with uninformed, model initialized thinking.
@@ 2.2 Iterative and Multi-Hop RAG Approaches
Chain-of-Thought (CoT) prompting (wei2022chain) encourages models to generate intermediate reasoning steps, effectively simulating an iterative thinking process. IRCoT (trivedi2022interleaving) demonstrated that interleaving retrieval and generation steps can significantly improve performance on multi-hop reasoning tasks. ITER-RETGEN (shao2023enhancing) proposed a framework where models can decide when to retrieve additional information based on their confidence levels. WebGPT (nakano2021webgpt) showed that models can be trained to browse the web iteratively to gather information for answering questions. ReAct (yao2023react) combined reasoning and acting in language models, enabling them to perform dynamic retrieval based on their reasoning traces. More recent work by (jiang2023active) introduced Self-RAG, which uses reflection tokens to control retrieval timing and assess the quality of retrieved passages,while Self-ask, proposed by (press2023measuringnarrowingcompositionalitygap), implements an autonomous question formulation mechanism during the reasoning process. FLARE (jiang2023activeretrievalaugmentedgeneration) incorporates adaptive retrieval when LLMs generate low-confidence tokens.
@@ 3.1 PPO
Proximal Policy Optimization (PPO) (schulman2017proximal) is an actor-critic reinforcement learning algorithm that has become the predominant method for RL fine-tuning of large language models (ouyang2022training). For language model fine-tuning, PPO maximizes the following objective:
@@ 3.2 GRPO
(shao2024deepseekmath) propose Group Relative Policy Optimization (GRPO), illustrated in Figure 2. GRPO eliminates the need for value function approximation by using the average reward of multiple sampled outputs as a baseline. For each question $q$, GRPO samples a group of $G$ outputs $\{o_{1},o_{2},\ldots,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and optimizes the following objective:

### v2

@@ 1 Introduction
Iterative RAG systems(Search-R1; R1-Searcher) are often optimized by Reinforcement Learning (RL)(PPO; GRPO), offering a principled approach to learn effective retrieval and reasoning strategies. RL-based RAG frameworks treat the retrieval and generation process as a sequential decision-making problem, where agents learn to search for information and generate responses to maximize cumulative rewards based on answer accuracy and efficiency metrics. The success of RL training heavily depends on the quality of the exploitation and the exploration efficiency during the learning process. Recent studies on entropy(wang2025beyond; cui2025entropy) show that entropy measurement is a good signal for this exploitation and exploration balance, which is important because the exploitation of retrieved information and exploration in the retrieval set control the whole reasoning process. Poor initial reasoning steps in exploration can lead to compounding errors throughout the iterative process.
@@
  • Analysis from an Entropy Perspective. We analyze the training dynamics of Group Relative Policy Optimization (GRPO)(GRPO) in iterative RAG from an entropy perspective and show that with lower entropy in the training phase, instead of insufficient exploration, our approach leads to more efficient exploration strategies focusing on the retrieval set, faster roads to the answer during RL training compared to traditional approaches that start with uninformed, model initialized thinking.
@@ 2.2 Iterative and Multi-Hop RAG Approaches
Chain-of-Thought (CoT) prompting (wei2022chain) encourages models to generate intermediate reasoning steps, effectively simulating an iterative thinking process. IRCoT (trivedi2022interleaving) demonstrated that interleaving retrieval and generation steps can significantly improve performance on multi-hop reasoning tasks. ITER-RETGEN (shao2023enhancing) proposed a framework where models can decide when to retrieve additional information based on their confidence levels. WebGPT (nakano2021webgpt) showed that models can be trained to browse the web iteratively to gather information for answering questions. ReAct (yao2023react) combined reasoning and acting in language models, enabling them to perform dynamic retrieval based on their reasoning traces. More recent work by (asai2024self) introduced Self-RAG, which uses reflection tokens to control retrieval timing and assess the quality of retrieved passages,while Self-ask, proposed by (press2022measuring), implements an autonomous question formulation mechanism during the reasoning process. FLARE (jiang2023active) incorporates adaptive retrieval when LLMs generate low-confidence tokens.
@@ 3.1 PPO
Proximal Policy Optimization (PPO) (PPO) is an actor-critic reinforcement learning algorithm that has become the predominant method for RL fine-tuning of large language models (ouyang2022training). For language model fine-tuning, PPO maximizes the following objective:
@@ 3.2 GRPO
(GRPO) propose Group Relative Policy Optimization (GRPO), illustrated in Figure 2. GRPO eliminates the need for value function approximation by using the average reward of multiple sampled outputs as a baseline. For each question $q$, GRPO samples a group of $G$ outputs $\{o_{1},o_{2},\ldots,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and optimizes the following objective: